A company makes pens. They sell each pen for $
5
.
Their revenue is represented by R = 5 x .
The cost to make the pens is $
1
each with a one time start up cost of $
4000
.
Their cost is represented by C = 1 x + 4000 .
a) Find the profit, P, (P = R - C) when the company sells 1000 pens.

Answer by ikleyn(53304)

.

They just explained everything to you in the post.   You only need to combine
their explanations in one simple formula,  and then calculate

        profit   P = R - C = 5x - (x + 4000) = 5*1000 - (1000+4000) = 5000 - 5000 = 0.

The profit is  0  (zero)  at the given conditions.

So,  x = 1000  is the break point in this problem.